# Improved bounds for randomized Schatten norm estimation of numerically low-rank matrices

### v2

@@ 3.2.4 Second-order analysis
The desired bound in the statement of theorem immediately follows by applying Lemmas 6, 9, 10, 13 and 14 to obtain
@@ 4 Numerical examples
We numerically illustrate the performance of our bounds and estimates on the variance of $\hat{\theta}_{2p}$. The numerical experiments have been performed in Matlab R2022b, and all the figures in this document can be reproduced using the code available at https://github.com/Alice94/schatten-norm. Recall that, without loss of generality, we can work with diagonal matrices. In the plots below, we estimate the exact variance of $\hat{\theta}_{2p}$ for a fixed matrix and a fixed value of $k$ by running $\hat{\theta}_{2p}$ for $10000$ times and computing the sample variance of the results; this corresponds to the blue line in the figures in this section. The red lines (first-order estimate) correspond to Theorem 3, the yellow lines (second-order estimate) correspond to Theorem 5, and the purple dotted lines correspond to the bound of Theorem 4. In all our examples, we keep the matrix size quite small ($100\times 100$) so that we are able to run the estimator many times to compute its empirical variance as precisely as possible.
@@
We consider the diagonal matrix $A\in\mathbb{R}^{100\times 100}$ with diagonal entries $0.8,0.8^{2},0.8^{3},\ldots,0.8^{100}$. Figure 3 illustrates the variance of $\hat{\theta}_{2p}$ for $p=4,6,8$ with values of $k$ ranging from $16$ to $500$. Note that the first-order estimate and the second-order estimates tend to underestimate the variance and get better for larger values of $k$, as expected.
Figure 3: Comparison of the variance of $\hat{\theta}_{2p}$, our bounds, and our estimates for the matrix from Example 15.
@@
In Figure 4, we illustrate the variance of $\hat{\theta}_{2p}$ for $p=3,5,7$ for the $100\times 100$ diagonal matrix with diagonal entries $1,1/4,1/9,1/16,\ldots,1/10000$.
Figure 4: Comparison of the variance of $\hat{\theta}_{2p}$, our bounds, and our estimates for the matrix from Example 16.
@@
The motivation for this work is to explore the behavior of $\hat{\theta}_{2p}$ for numerically rank matrices. However, even for the matrices with no eigenvalue decay, our bound from Theorem 4 improves the existing bound (2) from [6]. In Figure 5, we consider $A=I_{100}$, the identity matrix, and compare the empirical variance of $\hat{\theta}_{2p}$ with our estimates, our bounds, and the bound (2). Our bounds are not tight at all, especially for larger values of $p$, but are several orders of magnitude tighter than the bound (2).
Figure 5: Comparison of the variance of $\hat{\theta}_{2p}$, our bounds, our estimates, and bound (2) from [6] for the identity matrix.
@@ 5 Discussion
In this work, we analyzed the variance of $\hat{\theta}_{2p}$ given by Algorithm 1, which is an unbiased estimator for the $2p$-th power of the Schatten-$2p$ norm of a matrix $A$, assuming only a sketch $Y=A\Omega$ with a Gaussian matrix $\Omega$ is available. We provided a new bound – Theorem 4 – that improves the bound (2) from [6] for any matrix. Moreover, we focused on the matrices that have strong singular value decay, and proposed a first-order and second-order estimate on the variance. Numerically, we observe that these estimates are rather precise for moderate values of $p$ and $k$. For matrices with no singular value decay, the bound holds but the estimates do not really describe the behavior of $\hat{\theta}_{2p}$ since they are too optimistic.
The streaming setting we considered is applicable in scenarios where we need to extract information on the moments of a covariance matrix of a multivariate Gaussian distribution while only the samples drawn from the distribution is available. This is the original setting considered in [6]. Algorithm 1 was also suggested as a cheap Schatten-$2p$ norm estimator in the randomized numerical linear algebra review papers [8, 10]. We remark that, if the matrix $A$ can be accessed more than once via matrix-vector product, one should use other techniques to estimate Schatten norms; for instance, the ones based on Hutchinson trace estimator since they have smaller variance.

### v3

@@ 3.2.4 Second-order analysis
The desired bound in the statement of theorem immediately follows by applying Lemmas 9, 12, 13, 16 and 17 to obtain
@@ 4 Numerical examples
We numerically illustrate the performance of our bounds and estimates on the variance of $\hat{\theta}_{2p}$. The numerical experiments have been performed in Python, and all the figures in this document can be reproduced using the code available at https://github.com/Alice94/schatten-norm. Recall that, without loss of generality, we can work with diagonal matrices. In the plots below, we estimate the exact variance of $\hat{\theta}_{2p}$ for a fixed matrix and a fixed value of $k$ by running $\hat{\theta}_{2p}$ for $100000$ times if $k<100$ and for $3000$ times if $k\geq 100$, and computing the sample variance of the results. The sample variance corresponds to the blue line in the figures in this section. The orange lines (first-order estimate) correspond to Theorem 2, that is, we approximate the variance as $2p^{2}\|A\|_{4p}^{4p}/k$; the green lines (second-order estimate) correspond to Theorem 4, that is, we approximate the variance as
$\frac{2p^{2}\|A\|_{4p}^{4p}}{k}+\frac{p^{2}(p-1)^{2}}{k^{2}}\left(\|A\|_{4p}^{4p}+\frac{3}{2}\|A\|_{4}^{4}\|A\|_{4p-4}^{4p-4}-\frac{1}{2}\|A\|_{2p}^{4p}\right);$
the red dotted lines correspond to the bound in Theorem 3, and the purple dotted lines correspond to the bound (2). In all our examples, we keep the matrix size quite small ($100\times 100$) so that we are able to run the estimator many times to compute its empirical variance as precisely as possible.
@@
As a first example, we consider a matrix with singular values that exhibit exponential decay. We take $A$ to be the diagonal matrix $A\in\mathbb{R}^{100\times 100}$ with diagonal entries $0.8,0.8^{2},0.8^{3},\ldots,0.8^{100}$. Figure 3 illustrates the variance of $\hat{\theta}_{2p}$ for $p=4,6,8$ with values of $k$ ranging from $10$ to $1280$. Note that the first-order estimate and the second-order estimates tend to underestimate the variance and get better for larger values of $k$, as expected.
Figure 3: Comparison of the variance of $\hat{\theta}_{2p}$, our bounds, our estimates, and (2) for the matrix from Example 18.
@@
We now illustrate the behavior of our bounds for matrices whose singular values exhibit algebraic decay. In Figure 4 we consider the $100\times 100$ diagonal matrix with diagonal entries $1,1/4,1/9,1/16,\ldots,1/10000$, for $p=3,5,7$, and in Figure 5 we consider the $100\times 100$ diagonal matrix with diagonal entries $1,1/2^{4},1/3^{4},\ldots,1/100^{4}$ for $p=2,6,10$.
Figure 4: Comparison of the normalized variance of $\hat{\theta}_{2p}$, our bounds, our estimates, and (2) for the matrix from Example 19 with diagonal entries $1,1/4,1/9,\ldots,1/10000$.
Figure 5: Comparison of the variance of $\hat{\theta}_{2p}$, our bounds, our estimates, and (2) for the matrix from Example 19 with diagonal entries $1,1/2^{4},1/3^{4},\ldots,1/100^{4}$.
@@
The motivation for this work is to explore the behavior of $\hat{\theta}_{2p}$ for numerically low-rank matrices. However, even for the matrices with no eigenvalue decay, our bound from Theorem 3 improves the existing bound (2) from [10]. In Figure 6, we consider $A=I_{100}$, the identity matrix, and compare the empirical variance of $\hat{\theta}_{2p}$ with our estimates, our bounds, and the bound (2). Our bounds are not tight at all, especially for larger values of $p$, but are several orders of magnitude tighter than the bound (2).
Figure 6: Comparison of the variance of $\hat{\theta}_{2p}$, our bounds, our estimates, and (2) for the identity matrix.
Example 21.
We now consider a sequence of matrices of increasing size and same spectral density, and investigate the behavior of our bound and estimates. More precisely, we consider the diagonal matrix $A_{1}\in\mathbb{R}^{10\times 10}$ with diagonal elements $1,1/2^{4},\ldots,1/10^{4}$, and the matrices $A_{t}=\begin{bmatrix}A_{t-1}&0\\ 0&A_{t-1}\end{bmatrix}$ of increasing dimension, for $t=2,\ldots,10$, so that $A_{t}$ has size $10\cdot 2^{t-1}\times 10\cdot 2^{t-1}$. We fix $p=5$ and, for each matrix, we plot in Figure 7 the variance, the first-order and second-order estimates, our upper bound, and the bound (2), all of them divided by $\|A\|_{10}^{20}$. The number of samples is $k=100$. Note that the numerical rank increases linearly with the size of the matrix; this explains why our bounds and estimates are better for the matrices of smaller size.
Figure 7: Comparison of the variance of $\hat{\theta}_{2p}$, our bounds, our estimates, and (2) for the matrices from Example 21.
@@ 5 Discussion
In this work, we analyzed the variance of $\hat{\theta}_{2p}$ given by Algorithm 1, which is an unbiased estimator for the $2p$-th power of the Schatten-$2p$ norm of a matrix $A$, assuming only a sketch $Y=A\Omega$ with a Gaussian matrix $\Omega$ is available. We provided a new bound – Theorem 3 – that improves the bound (2) from [10] for any matrix. Moreover, we focused on the matrices that have strong singular value decay, and proposed a first-order and second-order estimate on the variance. Numerically, we observe that these estimates are rather precise for moderate values of $p$ and $k$. For matrices with no singular value decay, the bound holds but the estimates do not really describe the behavior of $\hat{\theta}_{2p}$ since they are too optimistic. In all cases, our upper bound on the variance was far tighter than the bound (2). After our paper was submitted, a preprint addressing the same problem of computing or approximating the variance of $\hat{\theta}_{2p}$ was posted to arXiv [8]; the upper bound in [8], however, explicitly depends on the size of the matrix, and is not tailored to the case of numerically low-rank matrices.
The sketching model we considered is applicable in scenarios where we need to extract information on the moments of a covariance matrix of a multivariate Gaussian distribution for which only some samples drawn from the distribution are available. This is the original setting considered in [10]. Algorithm 1 was also suggested as a cheap Schatten-$2p$ norm estimator in the randomized numerical linear algebra review papers [13, 15]. We remark that, if the matrix $A$ can be accessed more than once via matrix-vector product, one should use other techniques to estimate Schatten norms; for instance, the ones based on Hutchinson trace estimator since they have smaller variance.
Acknowledgments
The authors thank the anonymous referees for providing helpful comments that improved the quality of this work.